OpenStudy (anonymous):

Can anybody help me get the derivative? Question is: Find the tangent line to the folium of Descartes x^3 + y^3 = 6xy at the point (3,3). I put: x^3 + y^3 - 6xy = 0 Not sure how to get the derivative of -6xy

OpenStudy (amistre64):

product rule

OpenStudy (amistre64):

the constant rule allows states that a constant is useless so pull it out and save it for later

OpenStudy (amistre64):

the product rule for xy is: Dx(xy) = x'y + xy' ; where x' = 1 since dx/dx = 1

OpenStudy (amistre64):

bring back the -6 to get: -6(y + xy') = -6y -6xy'

OpenStudy (amistre64):

try to see the variable themselves as functions of some other "letter" x = x(t) and y = y(t) [x(t) y(t)]' = x'(t) y(t) + x(t) y'(t) ; by the product rule

OpenStudy (amistre64):

to wit: d (x(t)) d(x) ----- = --- dx dx x'(t) = dx/dx x'(t) = 1 ............................ d (y(t)) d(y) ----- = --- dx dx y'(t) = dy/dx y'(t) = y'

OpenStudy (anonymous):

ok im going to give this a shot. Thank you again. :)